OpenStudy (anonymous):

guys could you help me with this absolute value inequality |x-4| less or equal 1

OpenStudy (zzr0ck3r):

\(|a|\le b \iff -b\le a\le b\) so \(|x-4|\le1\iff -1\le x-4 \le 1\)

OpenStudy (zzr0ck3r):

solve for x

OpenStudy (anonymous):

omg i made i mistake the original thing is like this |4-x| less or equal 1